# Some models are useful, but for how long?: A decision theoretic approach to choosing when to refit large-scale prediction models

### v1

@@ 5 Discussion
As complex prediction models play ever more essential roles across nearly every sphere of inquiry, it has become more and more pertinent to answer the question: how much should one rely on an AI/ML model? By connecting this to a deep literature on portfolio optimization and the inference on predicted data paradigm [20], we suggest a new framework to evaluate the efficacy of an AI/ML model. We show that with this toolkit, one can create actionable decision rules based on your budget, the cost of model training, the cost of data, and the performance of the models, that can accurately determine the most efficient way to keep AI models up-to-date. Although we show this in the context of downstream statistical inference, the asset allocation framework has the potential to be extended to many other metrics used to evaluate AI/ML quality such as accuracy, fairness, and interpretability.
As researchers integrate AI technology across all levels, from domains where data collection is cheap and plentiful (such as web-based text or image data) to domains where data are siloed and expensive (such as healthcare), our framework provides new language to evaluate how to spend limited resources. This can be helpful in situations such as clinical trials, where there is the potential to intelligently replace human subjects with AI-approximated datapoints. This has the potential to save resources while ensuring that traditional efficacy and safety requirements are met. This can even be helpful in situations where the economics of data within a field is in flux.
Finally, we note that through this new connection to portfolio optimization theory, we are revisiting an old question of the value of synthetic data. In the last century, this was raised in statistics through the controversy over the two-phase sampling procedure [35, 10]. More recently, in the machine learning world, we have seen both that model training can be greatly improved by incorporating synthetic data [44, 36] and concerns that such procedures can lead to model collapse [4, 14]. In this paper, we study synthetic outcomes in terms of their ability to estimate outcomes for downstream statistical inference. By balancing the competing objectives of increasing accuracy and minimizing uncertainty, we observe that the value of synthetic data can vary wildy depending on one’s economic utility.
6 Limitations
@@
There are two important limitations. First, we assume that if a model were to be retrained in the future, one would get an MSE distribution that resembles the one during initial model training. This makes this procedure ill-suited when the signal-to-noise ratio in the data changes due to external factors influencing the “learnability” of the outcome. The second assumption is in the choice of how to forecast $\hat{W}$ and $\hat{D}$. We forecast both $\hat{W}$ and $\hat{D}$ using simple linear regression. In cases where more data is available, more complex forecasting methods can be used.

### v2

@@ 6 Discussion
As complex prediction models play ever more essential roles across nearly every sphere of inquiry, it has become more and more pertinent to answer the question: how much should one rely on outcomes predicted by an AI/ML model? By connecting this to a deep literature on portfolio optimization, we suggest a new framework for settings where the goal is statistical inference. After all, we cannot expect statistical uncertainty to vanish just because outcomes are predicted with sophisticated models. We show that with this toolkit, one can create actionable decision rules based on your budget, the cost of model training, the cost of data, and the performance of the models, which can accurately determine the most efficient way to keep relevant AI models up-to-date. We show this in the context of downstream statistical inference, but the asset allocation framework has the potential to be extended to many other metrics used to evaluate AI/ML quality such as accuracy, fairness, and interoperability.
Though this framework was designed with modern data science problems in mind, many of the topics connect to older discussions in statistics. Evaluating whether one’s resources are best spent collecting cheaper imputed or more expensive real outcomes has long been known in the sampling literature under the names of “two-phase sampling” and “double sampling” (Neyman, 1938; Hidiroglou and Särndal, 1998; Bose, 1943). However, much of this literature is developed in situations that require the imputation procedure or the downstream inference model to be well specified (Cochran, 1977; Fujii and van der Weide, 2019; Davidov and Haitovsky, 2000). However, such assumptions are difficult to verify when the purpose of many black-box procedures is that they model things that are difficult to interpret. Thus, newer approaches that will give unbiased estimates regardless of the specification of the true data generation or black-box procedure such as PPI++ are preferable.
@@
There are two important limitations of note. First, we assume that if a model were to be retrained in the future, one would get an MSE distribution that resembles the one during initial model training. This makes this procedure ill-suited when the signal-to-noise ratio in the data changes due to external factors influencing the “learnability” of the outcome. The second assumption is in the choice of how to forecast $\hat{W}$ and $\hat{D}$. We forecast both $\hat{W}$ and $\hat{D}$ using simple linear regression. In cases where more data is available, more complex forecasting methods can be used.
As researchers integrate AI technology across all levels, from domains where data collection is cheap and plentiful (such as web-based text or image data) to domains where data are siloed and expensive (such as healthcare), our framework provides a new language to evaluate how to spend limited resources. This can be helpful in situations such as clinical trials, where there is the potential to enrich human subjects with AI-approximated data points and save resources while meeting traditional efficacy and safety requirements.